Thursday, December 11: OWL Shift

From Hall A Wiki
Jump to: navigation, search

Go back to Daily Run Plan Daily Run Plans

  • General considerations
    • Kijun Park is the current Run Coordinator. Please contact directly via RC phone 757-876-1787
    • Follow the generic Instructions for shift takers
    • Shift leaders: Please make sure to fill out the beam time accounting form and write a "Shift Summary" log entry at the end of your shift regardless if there is beam or not.
    • In general, Hall technical staff should not be contacted by the shift crew - please call the RC first.
    • The target operator is required to supervise the cryo-target.
    • Make sure to use the DVCS CODA configuration for all runs (even when taking single arm HRS data).
    • Note that ion chamber calibrations have ALREADY been done for: carbon (single foil), carbon (five-foil), LH2, and BeO. See this log entry for the first three and this log entry for BeO. The limits on the ion chambers for these targets were set up for currents up to 20 uA.
  • Instructions for this shift
    • Beam centering (C hole target with Raster(2x2) on/off) to make sure a beam center info (Previously BMP-B [X:Y]=[-2.0:+2.5]) (Hopefully it takes less than 1hr)
      • simple check by event-rate on scaler with turning on/off raster with previous BPM-B values
    • Take a single run with beam 1, 2, 5, off, 10uA(or whatever maximum they can provide) for the BCM calib. [take around 1hr data taking]
    • Elastic calibration data (see white board for details, DVCS Calo is positioned at 6.138m + CH block was removed)
      • 1. setting#1: 3.18GeV/32.50(deg): 4-5hrs -> then, call Carlos 30min before
      • 2. setting#2: 2.95GeV/34.30(deg): 2-3hrs
      • 3. setting#3: 2.742GeV/36.20(deg): 1-2hrs